OpenStudy (anonymous):

Solve. Please show steps. lnx-ln(x-1)=1

OpenStudy (lalaly):

use \[\ln(\frac{a}{b})=lna-lnb\]

OpenStudy (lgbasallote):

then use \(\large \log_a b = c \rightarrow a^c = b\)

OpenStudy (lalaly):

easier way take e of both sides since its ln

OpenStudy (lgbasallote):

as the queen orders... \(\ln a = x \rightarrow e^x = a\)

OpenStudy (lalaly):

hehe try it @Chena804

OpenStudy (anonymous):

integration?

OpenStudy (anonymous):

\[\ln x-\ln (x-1)=1\] \[\ln (x \div(x-1))=1\] \[\left(\begin{matrix}x\\ x-1\end{matrix}\right)=e ^{1}\] \[x=(x-1)e ^{1}\] You can solve the rest. ;)